OpenStudy (anonymous):

At the outer edge of a rotating space habitat, 130 m from its center, the rotational acceleration is g. What is the rotational acceleration at a distance of 65 m from the center of the habitat?

OpenStudy (anonymous):

Is this the correct answer --- find the angular velocity by w^2R = g = 9.81 m/s^2 w^2 = 9.81/130 w = 0.2747 rad/s and at R = 65: a = w^2R = 0.2747^2*65 = 4.905 m/s^2 (= g/2)